OpenStudy (anonymous):

What are the diameter and radius of a circle with a circumference of 48π yd? A. d = 24 yd r = 12 yd B. d = 24 yd r = 48 yd C. d = 12 yd r = 24 yd D. d = 48 yd r = 24 yd please explain it beacause im so confused ithought the radius was 7.65 and the diameter was 15.3 BUT ITS NOT ITTTTT

jimthompson5910 (jim_thompson5910):

One circumference formula is C = pi*d where d is the diameter

jimthompson5910 (jim_thompson5910):

So if the given circumference is 48pi, then... C = pi*d 48pi = pi*d 48pi/pi = d ... Divide both sides by pi to isolate d. 48 = d d = 48

jimthompson5910 (jim_thompson5910):

Then you cut that diameter in half to get the radius r = d/2 r = 48/2 r = 24

OpenStudy (anonymous):

think you forgot it said \(48\pi\) and just thought if it as \(48\)

OpenStudy (anonymous):

OMG OMG THANK U SO MUCH YOU ARE THE BEST

jimthompson5910 (jim_thompson5910):

you're welcome